through (4,2) and (-4,-2)
Answer by xcentaur(357)

(4,2) and (-4,-2)


slope of the line (m) = difference in y-coordinates/difference in x-coordinates


m=2-(-2)/4-(-4)=4/8=1/2
Hence slope of the line=1/2


Taking any point,say(4,2) and using the point-slope form of straight line we get:


(y-2)=m(x-4)
y-2=1/2(x-4)
2y-4=x-4
2y=x
y=x/2


Equation of the line: [y=x/2]


Hope this helps,
Prabhat